# Representative Action Selection for Large Action-Space Meta-Bandits

### v1

@@ 4.1 Varying-dependence Actions
We study the effect of varying action dependence using a Gaussian process with a kernel. To control the degree of dependence, we use stationary RBF kernel and non-stationary Gibbs kernel (williams2006gaussian, ).
@@
where $l$ is a length-scale parameter and $l(a):=0.1+0.9\cdot\exp(-\|a\|^{2})$ is a location-dependent length-scale function. Both of them control the dependence between actions. But, unlike the stationary RBF kernels, the Gibbs kernel allows some regions of the action space to be highly correlated over large distances, and others to have short-range correlations. That is, the correlation depends not only on the distance between actions, but also on where $a,a^{\prime}$ are located. When $l(a)=l$ is a constant, the Gibbs kernel reduces to the RBF kernel.
Sampling Outcome Functions from a Kernel: We first construct the kernel matrix $\mathbf{K}$, where each entry is given by $\mathbf{K}_{a,a^{\prime}}=k(a,a^{\prime})$, for $a,a^{\prime}\in\mathcal{A}_{full}$, depending on the choice of kernel. We then sample a Gaussian vector (a Gaussian process function evaluated at finite input) $f\sim\mathcal{N}(0,\mathbf{K})$. Under either kernels defined in Equation (5), the variance of the function value $f(a)$ is one for all $a\in\mathcal{A}_{full}$.
Figure 2: Experiments with outcome functions sampled from RBF/Gibbs kernels in Equation (5). Left: Sampled outcome functions from Gibbs kernel over fixed 1000 grid points in $[0,2]$ (blue curves, right y-axis). The histogram (purple bars, left y-axis) shows action selection frequencies by Algorithm 1 with $K=5000$, favoring regions with rougher functions and edge points. Right: Performance on solving stochastic combinatorial optimization in Equation (6), selecting $K=5$ distinct actions from 15 grid points in [0,2], with outcome functions sampled from RBF kernel at varying length-scales. Solid curves and shaded regions represent the mean and $\pm$ standard deviation of the expected regret over 50 repetitions. Results are presented for our method (green), Thompson Sampling (yellow), and UCB (orange), with TS/UCB run for 3000 rounds per repetition.
To study the effect of varying dependence on the output of Algorithm 1, we consider a fixed action space consisting of 1000 grid points in the interval $[0,2]$, using Gibbs kernel in Equation (5) to sample outcome functions. To simplify computations, we marginalize a Gaussian process defined by the kernel over the grid. The left subplot of Figure 2 provides examples of sampled outcome functions (blue curves, with the y-axis on the right-hand side), which become smoother as the actions approach the left end of the interval—indicating stronger correlations among function values in that region. We run Algorithm 1 on this action space with $K=5000$ to select actions and record the frequency of each action being selected. The resulting histogram (purple bars, with the y-axis on the left) reflects the importance measure $q$, highlighting that Algorithm 1 tends to select more actions from regions where the outcome functions are rougher—i.e., where action outcomes are less correlated and their features $\Phi(a)$ are are farther apart. Another interesting aspect of this subplot is the two high bars at the edges. Recall that the actual action space consists of feature vectors $\Phi(a)$ for $a\in\mathcal{A}_{ful}$. For actions indexed closer to 0, their feature vectors become more densely packed compared to those indexed closer to 2, resulting in more correlated outcomes. The two actions at the edges, indexed by 0 and 2, correspond to the two farthest points in the actual feature space.
As an extreme case, suppose $\mathcal{A}_{full}$ is a set of independent actions. This corresponds to the canonical process with the action space given by the orthogonal basis in $\mathbb{R}^{n}$ where $n=|\mathcal{A}_{full}|$. To see this, we associate each action $a$ with a unit vector $e_{a}$, which is $1$ at the $a^{th}$ entry but is zero otherwise, and define the expected outcome as $\mu_{a}(\theta):=\langle e_{a},\theta\rangle$. With this construction, the collection $\{\mu_{a}\}_{a\in\mathcal{A}_{full}}$ consists of mutually independent random variables. Then, as discussed in Supplementary I, for any subset $\mathcal{A}\subset\mathcal{A}_{full}$, including the output of Algorithm 1, the regret suffers from the lower bound
$\mathbb{E}_{\theta}[\textnormal{Regret}]\geq\frac{\sqrt{\log|\mathcal{A}_{full}|}}{\sqrt{\pi\log 2}}-\sqrt{2\log|\mathcal{A}|}.$
4.2 Combinatorial Optimization under Uncertainty
@@
Returning to the objective stated in the Introduction, we aim to find a subset $\mathcal{A}$ of cardinality $K$ that minimizes the expected regret defined in Equation (1). Since the term $\mathbb{E}[\max_{a\in\mathcal{A}_{full}}\mu_{a}]$ is a constant independent of $\mathcal{A}$, this is equivalent to the following optimization problem:
@@
$\max_{\mathcal{A}\subseteq\mathcal{A}_{full}}\mathbb{E}\left[\max_{a\in\mathcal{A}}\mu_{a}\right]\quad\text{subject to }|\mathcal{A}|=K.$ (6)
@@
Our method is computationally efficient, avoiding the combinatorial complexity of the optimization problem in Equation (6). To assess the effectiveness in identifying near-optimal solutions, we compare it against Thompson Sampling (TS) and Upper Confidence Bound (UCB) on a simple example: We consider a fixed action space consisting of 15 grid points in $[0,2]$, using an RBF kernel in Equation (5) to sample outcome functions while varying the length-scale $l\in\{0.5,1,\dots,4\}$. In our method, we run Algorithm 1 until $K=5$ distinct actions are selected, where in each round, the optimal action is founded via an exhaustive search over the action space. In TS/UCB methods, we treat each $K$-tuple of actions as a super arm and adopt a bandit feedback setting: at each round, the decision-maker selects a super arm $\mathcal{A}$, observes the payoff $\max_{a\in\mathcal{A}}\mu_{a}$, and updates its policy accordingly, repeating this process for 3000 rounds. Thus, they are tailored to maximize the same objective. Since the payoff $\max_{a\in\mathcal{A}}\mu_{a}$ is unbounded, we assume Gaussian payoffs for both TS and UCB, with the algorithms summarized in Supplementary J, along with examples of selected subsets and sampled outcome functions. We evaluate expected regret of all methods over $10^{5}$ randomly sampled outcome functions.
@@
The right subplot of Figure 2 reports the expected regret of our method (green), TS (yellow) and UCB (orange), where solid curves and error shades indicate the mean $\pm$ one standard deviation of expected regret over 50 repetitions. As shown in this subplot, the expected regret decreases as the length-scale increases. This is because the length-scale $l$ of the RBF kernel controls the number of approximately independent actions in the process $\{\mu_{a}\}_{a\in\mathcal{A}_{full}}$: When actions $a$ and $a^{\prime}$ are far apart relative to the length-scale, the kernel value becomes very small, so the covariance between the function values $\mu_{a}$ and $\mu_{a^{\prime}}$ is also close to zero.
@@ Appendix A Connection to Bayesian Regret of Multi-Armed Bandits
Given a subset $\mathcal{A}\subseteq\mathcal{A}_{full}$, let $\mu^{*}(\mathcal{A},\theta):=\max_{a\in\mathcal{A}}\mu_{a}(\theta)$ be the highest expected outcome in a bandit instance $\theta$. We can equivalently rewrite the regret in Equation (1):
$\textnormal{Regret}(\theta)=\mu^{*}(\mathcal{A}_{full},\theta)-\mu^{*}(\mathcal{A},\theta).$
Bandit Regret: In the classic multi-armed bandit setting [7, 32], outcome of an arm $a\in\mathcal{A}_{full}$ in a bandit instance $\theta$ is generated i.i.d. according to an unknown but fixed law with expectation $\mu_{a}(\theta)$. A decision-maker would follow a policy $\pi$ that chooses the next arm based on the sequence of rounds and obtained outcomes. The bandit regret of policy $\pi$ in this bandit environment $\theta$ over $N$ rounds is defined as the expected difference in cumulative outcomes over $N$ rounds between always selecting the optimal action and following a policy $\pi$ to choose actions. Bayesian (bandit) regret [2, 32] is the expectation of this bandit regret, taken over the randomness in the distribution of bandit instances, and the action chosen by a policy:
$\begin{split}\textnormal{BanditRegret}(\theta):=&\mu^{*}(\mathcal{A}_{full},\theta)-\sum_{t=1}^{N}\mu_{A_{t}}(\theta),\\ \textnormal{BayesianRegret}:=&\mathbb{E}\left[\mu^{*}(\mathcal{A}_{full},\theta)-\sum_{t=1}^{N}\mu_{A_{t}}(\theta)\right],\end{split}$
where $A_{t}$ is the action chosen by policy $\pi$ in round $t$. Since we consider settings where only a subset of the full action space is accessible, the action chosen by the policy can only from the subset $\mathcal{A}\subseteq\mathcal{A}_{full}$. This quantifies the expected difference in cumulative outcomes over $N$ rounds between always selecting the optimal action from the full action set $\mathcal{A}_{full}$ and following policy $\pi$ to choose actions from the restricted subset $\mathcal{A}$.
Next, we establish the connection between the regret defined in this paper and the Bayesian (bandit) regret.
@@ A.1 Optimal policy
If $\pi$ is the optimal policy, it selects the best action in $\mathcal{A}$ in every round. Therefore, the regret of only having access to $\mathcal{A}$ instead of $\mathcal{A}_{full}$ for $N$ rounds, is determined by the optimal actions in $\mathcal{A}$ and $\mathcal{A}_{full}$:
@@ A.2 Thompson sampling policy
$=$ $\mathbb{E}\left[\sum_{t=1}^{N}\left(\mu^{*}(\mathcal{A})-\mu_{A_{t}}+\mu^{*}(\mathcal{A}_{full})-\mu^{*}(\mathcal{A})\right)\right]$
@@
where $|\mathcal{A}|$ denotes the cardinality, and $C>0$ is a constant. The inequality uses the Bayesian (bandit) regret of Thompson sampling (Theorem 36.1 in [32]). Note that with access to the full action space, the Bayesian regret is bounded by $C\sqrt{|\mathcal{A}_{full}|\cdot N\log N}$. However, this can be worse than restricting to a subset $\mathcal{A}$ when $\mathbb{E}_{\theta}[\textnormal{Regret}]$ is sufficiently small and the time horizon $N$ is finite. Intuitively speaking, using a well-chosen subset of representative actions instead of the full action space can reduce Bayesian regret in multi-armed bandits.
Appendix B Canonical Representation of a Gaussian Process
The expected outcome depends on both the chosen action $a\in\mathbb{R}^{n}$ and an observed context $g\in\mathbb{R}^{n}$, such that:
$\mu_{a}:=\langle a,g\rangle+c,\quad\;\forall\;a\in\mathcal{A}_{full},\;g\sim\mathcal{N}(0,\Sigma).$ (7)
We show how to transform it into the canonical Gaussian process. Note that
$g\equiv\Sigma^{1/2}\theta,\;\theta\sim\mathcal{N}(0,I),$
where $\equiv$ denotes equivalence in distribution. Let $\sigma_{j}$ denote the $j$-row of the matrix $\Sigma^{1/2}$. Thus, $g=(\langle\sigma_{j},\theta\rangle)_{j\leq n}$, and
$\mu_{a}=\langle a,g\rangle=\sum_{j\leq n}a_{j}g_{j}=\sum_{j\leq n}a_{j}\langle\sigma_{j},\theta\rangle=\left\langle\sum_{j\leq n}a_{j}\sigma_{j},\theta\right\rangle=\langle\Sigma^{1/2}a,\theta\rangle.$
Therefore, the setting in Equation (7) is equivalent to
$\mu_{a}=\langle a,\theta\rangle+c,\;\forall\;a\in\Sigma^{1/2}\mathcal{A}_{full},\;\theta\sim\mathcal{N}(0,I),$
where $\Sigma^{1/2}\mathcal{A}_{full}$ is the projected set of $\mathcal{A}_{full}$ by the matrix $\Sigma^{1/2}$.
Appendix C Convex hull dominated by extreme points
Lemma 9.
Given a subset $\mathcal{A}\subset\mathcal{A}_{full}$. If $\mathcal{A}$ includes all extreme points of $\operatorname*{Conv}(\mathcal{A}_{full})$, i.e., $\mathcal{A}_{full}\subseteq\operatorname*{Conv}(\mathcal{A})$, then it always holds
$\max_{a\in\mathcal{A}_{full}}\mu_{a}-\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}=0\;\;\forall\theta\in\mathbb{R}^{n}.$
Proof.
Let $\mathcal{A}=\{a_{1},a_{2},\dots,a_{K}\}$. Since $\mathcal{A}_{full}\subseteq\operatorname*{Conv}(\mathcal{A})$, for any $a\in\mathcal{A}_{full}$, there exist non-negative numbers $\lambda_{1},\dots,\lambda_{K}$ that sum up to $1$ and such that $a=\lambda_{1}a_{1}+\dots+\lambda_{K}a_{K}$. Thus, by Equation (2), for any $\theta\in\mathbb{R}^{n}$
$\mu_{a}=\langle a,\theta\rangle=\left\langle\sum_{i\in 1:K}\lambda_{i}a_{i},\theta\right\rangle=\sum_{i\in 1:K}\lambda_{i}\langle a_{i},\theta\rangle\leq\sum_{i\in 1:K}\lambda_{i}\max_{a^{\prime}\in\mathcal{A}}\langle a^{\prime},\theta\rangle=\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}.$
It yields
$\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}}\leq\max_{a\in\mathcal{A}_{full}}\mu_{a}\leq\max_{a^{\prime}\in\mathcal{A}}\mu_{a^{\prime}},$
where the first inequality uses $\mathcal{A}\subset\mathcal{A}_{full}$. Thus, the two quantities are equal. ∎
Appendix D Proof of Lemma 3

### v2

@@ 4.1 Varying-dependence Actions
Assuming the outcome function is generated from a kernel inherently implies that the expected outcomes of actions are correlated. To study the effect of varying correlation levels among actions, we conduct the following experiments:
Sampling Outcome Functions from a Kernel.
We use the stationary RBF kernel (Williams and Rasmussen, 2006), and provide additional illustrations using the non-stationary Gibbs kernel in Supplementary:
@@
where $l$ is a length-scale parameter that control the dependence between actions. By (Kanagawa et al., 2018, Theorem 4.12), we sample functions from a RKHS by first constructing the kernel matrix $\mathbf{K}$ with entries $\mathbf{K}_{a,a^{\prime}}=k(a,a^{\prime})$ for $a,a^{\prime}\in\mathcal{A}_{\operatorname*{full}}$, and then drawing $f\sim\mathcal{N}(0,\mathbf{K})$.
Figure 1: Performance comparison on stochastic combinatorial optimization in Equation (6), selecting $K=5$ distinct actions from 15 grid points, with outcome functions $f(a):=\mu_{a}$ sampled from an RBF kernel at varying length-scales. Left: Expected regret for our method (green), Thompson Sampling (yellow), UCB (orange), and Successive Halving (SH, blue), averaged over 50 repetitions at each length-scale. TS/UCB are run for 3000 rounds; SH is given a budget of 37,000 pulls. Right: SH’s expected regret (blue) and pull counts (yellow bars) per round. Our method’s performance (green) is shown for comparison. SH requires nearly 10,000 pulls—about 3× the number of super arms—to match our method, making it infeasible for large $\mathcal{A}_{\operatorname*{full}}$.
@@
Returning to the objective stated in the Introduction, we aim to find a subset $\mathcal{A}$ of cardinality $K$ that minimizes the expected regret defined in Equation (1). Since the term $\mathbb{E}[\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}]$ is a constant independent of $\mathcal{A}$, this is equivalent to the following optimization problem:
@@
$\max_{\mathcal{A}\subseteq\mathcal{A}_{\operatorname*{full}}}\mathbb{E}\left[\max_{a\in\mathcal{A}}\mu_{a}\right]\quad\text{subject to }|\mathcal{A}|=K.$ (6)
@@
Our method is computationally efficient, avoiding the combinatorial complexity of the optimization problem in Equation (6). To assess effectiveness in identifying near-optimal solutions, we compare our method against Thompson Sampling (TS) and Upper Confidence Bound (UCB) in a simple example, using Successive Halving (SH) as a reference for approximating the optimal subset—though SH becomes impractical when the full action space is large.
@@
We consider a fixed action space consisting of 15 grid points in [0,2][0,2][ 0 , 2 ], using an RBF kernel in Equation (5) to sample outcome functions while varying the length-scale $l\in\{0.5,1,\dots,4\}$. In our method, we run Algorithm 1 until $K=5$ distinct actions are selected, where the payoff of each action in a round is given by $\mu_{a}$. We use exhaustive search over the action space to find the optimal action $\arg\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}$. For other methods, we treat each $K$-tuple of actions as a super arm, and the payoff of each super arm in a round $\mathcal{A}$ is given by $\max_{a\in\mathcal{A}}\mu_{a}$. Thus, these methods are tailored to find the best super arm that maximizes $\mathbb{E}\left[\max_{a\in\mathcal{A}}\mu_{a}\right]$, aligning with the same objective. In TS/UCB methods, we adopt a bandit feedback setting: at each round, the decision-maker selects a super arm $\mathcal{A}$, observes the payoff $\max_{a\in\mathcal{A}}\mu_{a}$, and updates its policy accordingly, repeating this process for 3000 rounds, chosen to roughly match the number of super arms ($N=3003$). Since the payoff $\max_{a\in\mathcal{A}}\mu_{a}$ is unbounded, we assume Gaussian payoffs for both TS and UCB. The algorithms are summarized in Supplementary, with the prior of TS set to $\mathcal{N}(0,1)$. In the SH method, all super arms are evaluated using a fixed budget of arm pulls (we use the minimum budget required in Karnin et al. (2013), $N\log_{2}N\approx 37{,}000$) over a few rounds. In each round, the number of remaining super arms is halved. This process continues until only the best-performing super arm remains or the budget is exhausted. We evaluate expected regret of all methods over $10^{5}$ randomly sampled outcome functions.
@@
The left subplot of Figure 1 reports the expected regret of our method (green), TS (yellow), UCB (orange) and SH (blue), where solid curves and error shades indicate the mean $\pm$ one standard deviation of expected regret over 50 repetitions. As shown in this subplot, the expected regret decreases as the length-scale increases. This is because the length-scale $l$ of the RBF kernel controls the number of approximately independent actions in the process $\{\mu_{a}\}_{a\in\mathcal{A}_{\operatorname*{full}}}$: When actions $a$ and $a^{\prime}$ are far apart relative to the length-scale, the kernel value becomes very small, so the covariance between the function values $\mu_{a}$ and $\mu_{a^{\prime}}$ is close to zero.
Regarding computational complexity: in a given bandit instance $\theta$, identifying the optimal action $a^{*}(\theta):=\arg\max_{a\in\mathcal{A}_{\operatorname*{full}}}\mu_{a}(\theta)$ is a well-studied problem in both best-arm identification and GP optimization. However, if computing $a^{*}(\theta)$ is not computationally efficient, then identifying the optimal super arm, as in SH—whose search space grows combinatorially—is infeasible. Next, in the right subplot of Figure 1, we examine sample complexity. We run our method (green) and Successive Halving (SH, blue) for 50 additional repetitions, using a fixed length-scale $l=1$. Since SH proceeds in rounds, we track the current best super arm and the total number of super arm pulls so far after each round. The expected regret of the current best super arm, is evaluated over another $10^{5}$ randomly sampled outcome functions. The blue solid curves and shades for SH represent the mean $\pm$ one standard deviation of expected regret over 50 repetitions over each SH round. The yellow bars (corresponding to the right y-axis) show the number of super arm pulls by SH at the end of each round. For comparison, the green curve and shaded region represent the mean $\pm$ one standard deviation of expected regret for our method over the same 50 repetitions. As shown, SH’s performance roughly matches ours after three rounds, requiring 9005 super arm pulls—approximately three times the number of super arms. If the action space $|\mathcal{A}_{\operatorname*{full}}|$ is already large, running SH is intractable.
4.2 Conclusion and Future Work
We proposed a framework for selecting a subset of correlated actions, modeling payoff correlations with a Gaussian process and extending to the sub-Gaussian case. A simple algorithm was introduced and shown to effectively identify near-optimal subsets. A key direction for future work is developing a stopping criterion. When the subset size is flexible but sampling new bandit instances is costly, ideas from species discovery Roswell et al. (2021) may be useful—treating the discovery of a new action as analogous to discovering a new species. One could use Turing’s formula to estimate the probability of finding a new action, or frame the process as a one-armed bandit that stops when the discovery probability becomes sufficiently low.
@@ Appendix A Connection to Bayesian Regret of Multi-Armed Bandits
Given a subset $\mathcal{A}\subseteq\mathcal{A}_{\operatorname*{full}}$, let $\mu^{*}(\mathcal{A},\theta):=\max_{a\in\mathcal{A}}\mu_{a}(\theta)$ be the highest expected outcome in a bandit instance $\theta$. We can equivalently rewrite the regret in Equation (1):
$\textnormal{Regret}(\theta)=\mu^{*}(\mathcal{A}_{\operatorname*{full}},\theta)-\mu^{*}(\mathcal{A},\theta).$
Bandit Regret: In the classic multi-armed bandit setting [Auer et al., 2002a, Lattimore and Szepesvári, 2020], outcome of an arm $a\in\mathcal{A}_{\operatorname*{full}}$ in a bandit instance $\theta$ is generated i.i.d. according to an unknown but fixed law with expectation $\mu_{a}(\theta)$. A decision-maker would follow a policy $\pi$ that chooses the next arm based on the sequence of rounds and obtained outcomes. The bandit regret of policy $\pi$ in this bandit environment $\theta$ over $N$ rounds is defined as the expected difference in cumulative outcomes over $N$ rounds between always selecting the optimal action and following a policy $\pi$ to choose actions. Bayesian bandit regret [Agrawal and Goyal, 2012, Lattimore and Szepesvári, 2020] is the expectation of this bandit regret, taken over the randomness in the distribution of bandit instances, and the action chosen by the policy:
$\begin{split}\textnormal{BanditRegret}(\theta):=&\mu^{*}(\mathcal{A}_{\operatorname*{full}},\theta)-\sum_{t=1}^{N}\mu_{A_{t}}(\theta),\\ \textnormal{BayesianBanditRegret}:=&\mathbb{E}\!\left[\mu^{*}(\mathcal{A}_{\operatorname*{full}},\theta)-\sum_{t=1}^{N}\mu_{A_{t}}(\theta)\right],\end{split}$
where $A_{t}$ is the action chosen by policy $\pi$ in round $t$. Since we consider settings where only a subset of the full action space is accessible, the action chosen by the policy can only from the subset $\mathcal{A}\subseteq\mathcal{A}_{\operatorname*{full}}$. This quantifies the expected difference in cumulative outcomes over $N$ rounds between always selecting the optimal action from the full action set $\mathcal{A}_{\operatorname*{full}}$ and following policy $\pi$ to choose actions from the restricted subset $\mathcal{A}$.
Next, we establish the connection between the regret defined in this paper and the Bayesian bandit regret.
@@ A.1 Optimal policy
If $\pi$ is the optimal policy, it selects the best action in $\mathcal{A}$ in every round. Therefore, the regret of only having access to $\mathcal{A}$ instead of $\mathcal{A}_{\operatorname*{full}}$ for $N$ rounds, is determined by the optimal actions in $\mathcal{A}$ and $\mathcal{A}_{\operatorname*{full}}$:
@@ A.2 Thompson sampling policy
$=$ $\mathbb{E}\left[\sum_{t=1}^{N}\left(\mu^{*}(\mathcal{A})-\mu_{A_{t}}+\mu^{*}(\mathcal{A}_{\operatorname*{full}})-\mu^{*}(\mathcal{A})\right)\right]$
@@
where $|\mathcal{A}|$ denotes the cardinality, and $C>0$ is a constant. The inequality uses the Bayesian (bandit) regret of Thompson sampling (Theorem 36.1 in Lattimore and Szepesvári [2020]). Note that with access to the full action space, the Bayesian regret is bounded by $C\sqrt{|\mathcal{A}_{\operatorname*{full}}|\cdot N\log N}$. However, this can be worse than restricting to a subset $\mathcal{A}$ when $\mathbb{E}_{\theta}[\textnormal{Regret}]$ is sufficiently small and the time horizon $N$ is finite. Intuitively speaking, using a well-chosen subset of representative actions instead of the full action space can reduce Bayesian regret in multi-armed bandits.
Appendix B Proof of Lemma 4